What is the mass, in g, of 2.39 ×1015 platinum (Pt) atoms? Enter your answer in scientific notation.
Answer #1

Number of atoms of platinum= 2.39×10^15

Since 1 mole of atom= 6.022× 10^23 atoms

Let's suppose x mole of atom contains 2.39×10^15 atoms

Then x mole= (2.39×10^15/6.022×10^23)

=> 3.96×10^-9 mol

Molar mass of platinum= 195.084 g/mol

Since mass of 1 mole of platinum= 195.084 g/mol

Therefore mass of 3.96×10^-9 mol of platinum=

195.084× 3.96×10^-9= 772.53×10^-9 gram

Therefore mass of 2.39×10^15 atoms of platinum in gram= 772.53×10^-9